OpenStudy (anonymous):

Find the length of the curve: y = integral of x to 1, sqrt (sqrt(t) -1)dt where 1≤x≤16 How do I get this in terms that I can understand?

OpenStudy (amistre64):

use the draw button to try to write that out better, hard to read the information

OpenStudy (amistre64):

or if you can take a screenshot or pic of the problem that would be good as well

OpenStudy (anonymous):

\[\int\limits_{1}^{x}\sqrt{\sqrt{t}-1}dt\] where 1≤x≤16 is the original equation

OpenStudy (amistre64):

interesting ....

OpenStudy (amistre64):

\[ds=\sqrt{(x')^2+(y')^2}\] the length of a curve is just:\[\int ds\] you say:\[y=\int\limits_{1}^{x}\sqrt{\sqrt{t}-1}dt\] \[y'=\sqrt{\sqrt{x}-1}(x')-\sqrt{\sqrt{1}-1}(1')\]

OpenStudy (amistre64):

y'^2 = sqrt(x)-1 x'^2 = 1 ds = sqrt(1+sqrt(x)-1), from 1 to 16

OpenStudy (amistre64):

\[L=\int_{1}^{16}\sqrt{x^{1/2}}~dx\] \[L=\int_{1}^{16}x^{1/4}~dx\]

OpenStudy (anonymous):

Oh, I got it now! So the derivative of a derivative is plugging in the points of the integral and subtracting both ends?

OpenStudy (amistre64):

in general:\[F(x)=\int_{\alpha}^{\beta}f(t)~dt\] \[F(x)=F(\beta)-F(\alpha)\] taking the derivative invokes the chain rule \[\frac{d}{dx}F(x)=f(\beta)~\beta'-~f(\alpha)~\alpha'\]

OpenStudy (amistre64):

in this case: 1' = 0, and x' = 1 leaving us the original inset expression in terms of x

OpenStudy (anonymous):

Thank you so much!

OpenStudy (amistre64):

youre welcome